OpenStudy (anonymous):

decide which of the following have sums that are less than one a 1/3+2/5 b 7/8+1/2 c 5/6+2/3 d 1/7+3/9

OpenStudy (anonymous):

i know its not b or c

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@adamaero @IAmSinged

OpenStudy (anonymous):

its b

OpenStudy (adamaero):

I thought you got the hang of these already?

OpenStudy (anonymous):

haha kinds sorry i just really genuinely suck at these, sorry

OpenStudy (anonymous):

its okay

sammixboo (sammixboo):

Yes, we know it isn't B or C. B 7/8 is more than one-half and we have one-half, that means the fractions would equal over 1 C 5/6 is over one-half and so is 2/3, so therefore it would be over 1 Now let's see if A or D is our answer Choice A \(\rm \color{#20bd23}{\dfrac{1}{3}+\dfrac{2}{5}~=~\dfrac{5}{15}+\dfrac{6}{15}}\) All you need to do is add the numerators (the top numbers of the fractions). If the sum of the numerator is 15+ then this is not your answer

OpenStudy (anonymous):

okay ty!

OpenStudy (anonymous):

d probably